# Efficient Reasoning

### v2

@@ 1 Introduction
Masked diffusion language models (MDLM) have recently advanced rapidly, with pre-trained models such as LLaDA-8B (nie2025large) and Dream-7B (ye2025dream7b) scaling discrete diffusion and demonstrating strong performance across language, math, and code. Inference in these models is semi-autoregressive: the sequence is partitioned into contiguous blocks $B_{1},\dots,B_{T}$ that are generated left-to-right (autoregressive across blocks), while within each block decoding proceeds by diffusion-style denoising steps $s=1,\dots,S$ that can unmask any subset of still-masked positions in a non-left-to-right order. At each step, a mask predictor proposes token distributions and a fixed per‑step quota of token positions is filled with topk by confidence or randomly, while the remaining positions stay masked for later steps.
To push throughput beyond fixed number of tokens per step, recent works like Fast-dLLM (wu2025fastdllm) enables parallel decoding by unmasking all tokens above a static global confidence threshold. However, our empirical observations show that confidence is highly non-uniform: it varies across blocks and denoising steps, yet confidence trajectories are remarkably stable across inputs within the same dataset (or, the same task). The first observation calls for adaptivity over the course of generation; the second indicates that a small amount of computation can capture a task-level “confidence signature.” Together, they reveal substantial space for more efficient parallel decoding in MDLM than static schedules can offer.
@@ 2 Observation
Prior work wu2025fastdllm assumes confidence follows a generalizable pattern and applies a static threshold across tasks. In practice, as shown in Figure 1, benchmarks as different as GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation) exhibit distinct signatures. This task-dependent variability undermines any “one-size-fits-all” thresholding strategy, highlighting the need for adaptation at inference time.
@@ 4 Evaluation
We evaluate One-Shot Dynamic Thresholding (OSDT) on three benchmarks: GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation). OSDT calibrates a confidence profile on the first sequence of a dataset and applies dynamically adjusted thresholds to subsequent inputs. We study both its hyperparameter behavior and its performance against Fast-dLLM (wu2025fastdllm). All experiments use batch size 1 on a single NVIDIA H100 GPU.
@@ 5 Related Work
Recent advances in masked diffusion language models (DLMs) demonstrate their potential as an alternative to autoregressive generation. LLaDA (nie2025large) scales discrete diffusion to 8B parameters with block-wise decoding and low-confidence remasking, while Dream (ye2025dream7b) improves training via AR initialization and context-adaptive token-level noise rescheduling. Both rely on fixed-step schedules that unmask tokens one by one, limiting efficiency.
Fast-dLLM (wu2025fastdllm) addresses this by formally proving when greedy parallel decoding with product-of-marginals is equivalent to sequential decoding in high-confidence regimes, enabling safe parallel generation. It then proposes a confidence-aware parallel decoding rule that unmasks all tokens above a global static cutoff, and introduces prefix and dual (prefix+suffix) KV-Cache designs to improve throughput. However, its thresholding remains task-agnostic and static.

### v3

@@ 1 Introduction
Masked diffusion language models (MDLM) have recently advanced rapidly, with pre-trained models such as LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025) scaling discrete diffusion and demonstrating strong performance across language, math, and code. Inference in these models is semi-autoregressive: the sequence is partitioned into contiguous blocks $B_{1},\dots,B_{T}$ that are generated left-to-right (autoregressive across blocks), while within each block decoding proceeds by diffusion-style denoising steps $s=1,\dots,S$ that can unmask any subset of still-masked positions in a non-left-to-right order. At each step, a mask predictor proposes token distributions and a fixed per‑step quota of token positions is filled with topk by confidence or randomly, while the remaining positions stay masked for later steps.
To push throughput beyond fixed number of tokens per step, recent works like Fast-dLLM (Wu et al., 2025) enables parallel decoding by unmasking all tokens above a static global confidence threshold. However, our empirical observations show that confidence is highly non-uniform: it varies across blocks and denoising steps, yet confidence trajectories are remarkably stable across inputs within the same dataset (or, the same task). The first observation calls for adaptivity over the course of generation; the second indicates that a small amount of computation can capture a task-level “confidence signature.” Together, they reveal substantial space for more efficient parallel decoding in MDLM than static schedules can offer.
@@ 2 Observation
Prior work Wu et al. (2025) assumes confidence follows a generalizable pattern and applies a static threshold across tasks. In practice, as shown in Figure 1, benchmarks as different as GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation) exhibit distinct signatures. This task-dependent variability undermines any “one-size-fits-all” thresholding strategy, highlighting the need for adaptation at inference time.
@@ 4 Evaluation
We evaluate One-Shot Dynamic Thresholding (OSDT) on three benchmarks: GPQA (expert-level Q&A), GSM8K (grade-school math), and HumanEval (code generation). OSDT calibrates a confidence profile on the first sequence of a dataset and applies dynamically adjusted thresholds to subsequent inputs. We study both its hyperparameter behavior and its performance against Fast-dLLM (Wu et al., 2025). All experiments use batch size 1 on a single NVIDIA H100 GPU.
@@ 5 Related Work
Recent advances in masked diffusion language models (DLMs) demonstrate their potential as an alternative to autoregressive generation. LLaDA (Nie et al., 2025) scales discrete diffusion to 8B parameters with block-wise decoding and low-confidence remasking, while Dream (Ye et al., 2025) improves training via AR initialization and context-adaptive token-level noise rescheduling. Both rely on fixed-step schedules that unmask tokens one by one, limiting efficiency.
Fast-dLLM (Wu et al., 2025) addresses this by formally proving when greedy parallel decoding with product-of-marginals is equivalent to sequential decoding in high-confidence regimes, enabling safe parallel generation. It then proposes a confidence-aware parallel decoding rule that unmasks all tokens above a global static cutoff, and introduces prefix and dual (prefix+suffix) KV-Cache designs to improve throughput. However, its thresholding remains task-agnostic and static.